OpenStudy (anonymous):

d=vi+ 1/2at^2 solve for a vf^2=vi^2+2ad solve for vi and a

OpenStudy (anonymous):

2(d-vi)/t^2

OpenStudy (anonymous):

(vf^2-vi^2)/2d=a \[\sqrt{(vf^2-2ad)}\]=vi

OpenStudy (anonymous):

what about the first one

OpenStudy (anonymous):

how would i solve for a

OpenStudy (anonymous):

@protosbg